OpenStudy (anonymous):

How to find the first five coefficients of the power series 10/(1+49x^2)?

OpenStudy (anonymous):

think if it as a geometric series

OpenStudy (anonymous):

I understand that part. I don't know how to find the coefficients reliably.

OpenStudy (anonymous):

treat is as \[\frac{1}{1+(9x)^2}\] \[=1-(9x)+(9x)^2-(9x)^3+...\]

OpenStudy (anonymous):

Do you mean \[1/(1 - (-7x)^2) ? \]

OpenStudy (anonymous):

where r = (-7x)? because 9*9 is 81 not 49